the cost of a gym membership includes the initial fee and a monthly fee. Find the initial and monthly fees if the total cost for 5 months is $225 and the total cost for 10 months is $375.
Answer by macston(5194)

.
M=monthly fee; F=initial fee
.
Total cost=initial fee+(number of months)(monthly fee)
.
$225=F+5M
$225-5M=F
.
$375=F+10M
$375=($225-5M)+10M
$375=$225+5M
$150=5M
$30=M
ANSWER 1: The monthly fee is $30.
.
$225=F+5M
$225=F+5($30)
$225=F+$150
$75=F
ANSWER 2: The initial fee is $75
.
CHECK:
$375=F+10M
$375=$75+10($30)
$375=$75+$300
$375=$375